OpenStudy (anonymous):

5x-12=2x+6 sole equation?

OpenStudy (anonymous):

like terms so 5x-2x=6+12

OpenStudy (anonymous):

what do u get

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

no x does not equal 2 it equals 6

OpenStudy (anonymous):

3x=18 x=6

OpenStudy (anonymous):

mistake in + or - it is my fault

OpenStudy (anonymous):

5x - 12 = 2x + 6 Add 12 to both sides. 5x = 2x + 18 Subtract 2x from both sides. 3x = 18 Divide both sides by 3. x = 6